# Green Edge AI: A Contemporary Survey

### v1

@@ IV-B Model and Gradient Compression
In each training iteration, the model updates to be uploaded by EUDs take the form of model weights or model gradients. Compressing these model updates can reduce communication energy consumption, as fewer bits need to be transmitted. In addition, locally training compressed models can effectively minimize on-device computation energy consumption, as it can avoid substantial FLOPs. Typical compression approaches such as quantization, sparsification, and knowledge distillation are often employed to achieve energy-efficient FEEL.
@@ IV-B1 Quantization
Quantization employs low bitwidths to represent DNN model parameters. This technique can also be applied to model gradients prior to transmission. Given that quantization is an irreversible process, it is crucial to balance achieving maximum EE with maintaining satisfactory learning performance.
@@ IV-B3 Knowledge Distillation
A well-known KD-based FL algorithm called FedMD [134] addresses this issue by allowing EUDs to have heterogeneous model architectures based on their resource constraints. To facilitate cooperative training, logits computed from a public dataset using local models are uploaded to and aggregated at the MEC server. These aggregated logits are then redistributed to EUDs for local training using TL and KD. Since logits are of low dimension, the communication overhead of FedMD is significantly reduced compared to FedAvg. Empirical results in [135] provide further insights into key parameters of KD-based FEEL, such as the size and distribution of the public dataset, as well as the logit aggregation strategy. Another approach, FedKD [136], proposes maintaining a teacher and a student model at each EUD that can learn from each other, with only the approximated student model gradients, which have much smaller sizes, being uploaded. Complexity analysis demonstrates that FedKD significantly reduces communication costs with a marginal increase in local computation energy consumption. KD can also be implemented at the MEC server to simplify computation at EUDs. For example, in Mix2FLD [137], each EUD uploads a set of mixed-up data samples along with the logits. The logit-to-model conversion is performed by running a KD algorithm at the MEC server, after which the updated global model is downloaded by the EUDs. Mix2FLD demonstrates faster convergence and higher model accuracy, implying much lower energy consumption at the EUDs.
@@ IV-C1 Local Training Adaptation
Since most computations in FEEL are disseminated to EUDs, the approach to local training in each iteration has direct impacts on learning performance, as well as time and energy efficiency. Intuitively, better training of local models in each iteration reduces the required number of iterations to achieve a learning performance target [114, 138], but it increases on-device processing and computation energy consumption per iteration. However, a smaller number of iterations is advantageous for saving communication energy throughout the training process. Therefore, local training in FEEL should be adapted to find the optimal balance.
In [139], local training adaptation was first studied for FEEL, where a convergence bound was derived considering arbitrary local training epochs in each global iteration. This bound was utilized to determine the number of local epochs that minimize the loss function while adhering to resource constraints, such as energy and monetary costs. The batch size of local SGD also plays a significant role, as a larger batch size reduces gradient variance but increases computation overhead [140]. To conserve battery energy by minimizing device idling, an efficient algorithm was developed in [141] to dynamically determine batch sizes and learning rates for heterogeneous EUDs. Additionally, scaling the on-device processors to appropriate speeds can improve the EE of local training. For instance, in [142], the total system cost of FEEL defined as a weighted sum of training time and energy consumption was minimized using an experience-driven central processing unit (CPU)-cycle frequency control algorithm. Compared to using a static CPU-cycle frequency, this approach achieves over $4\%$ energy savings with a substantial reduction in training time.
@@ IV-C2 Joint Computation-Communication Control
Recent studies have also incorporated state-of-the-art computing architectures [146] and air interface technologies [147]. In [146], each EUD is assumed to be equipped with a CPU-GPU heterogeneous computing platform. Consequently, the joint optimization of communication resource management, CPU-GPU workload partitioning, and processor frequency scaling was performed to minimize total energy consumption. Experimental results demonstrated the effectiveness of joint resource management in improving EE compared to separate management of heterogeneous resources. Attention was also given to the intelligent reflecting surface (IRS) technology, which utilizes passive elements such as diodes and phase shifters to customize wireless environments in an energy-efficient manner. In [147], an IRS was proposed to be deployed in FEEL systems to provide favorable wireless channels, with the phase shifts at the IRS and the local CPU speed being judiciously designed. Simulation results showcased speedy convergence and significant energy savings, especially with a large-scale IRS.
@@ IV-C3 Device Selection
In [150], a device selection problem was tackled by a hybrid branch-and-bound algorithm in each training iteration to minimize the energy consumption of participating EUDs in a TDMA-based FEEL system. Additionally, the trade-off between maximizing the number of selected EUDs and minimizing their total energy consumption was investigated in [151], leading to the development of an energy- and latency-aware device selection algorithm. Both [150] and [151] demonstrate the effectiveness of device selection in achieving energy-efficient FEEL.
The aforementioned studies focus on optimizing device selection independently in each training iteration. However, all decisions made throughout the training process can have a significant impact on the learning outcome. This motivates the design of device selection strategies from a long-term perspective [152, 153, 154, 155]. In particular, recognizing that the number of participating EUDs in later training rounds has a more profound impact on learning performance, device selection and bandwidth allocation were jointly optimized in [152] using Lyapunov optimization, subject to energy constraints at each EUD. Experimental results demonstrate the benefits of this approach over methods that myopically minimize energy consumption in each round. This work was further extended in [153] to address the device heterogeneity issue by exploiting sequential transmission of local model updates, where early available local model updates can be uploaded to the MEC server using more bandwidth. Unlike [152] and [153], which use the average number of participating EUDs as a surrogate for learning performance, [154] developed an online data-driven method to estimate learning performance using a validation dataset. Based on this estimation, the selected devices and resource management were jointly optimized to minimize the system energy consumption while respecting latency and accuracy requirements. In [155], a lightweight AutoFL framework was developed based on reinforcement learning to select the best EUDs in each traininng round, achieving 5.2×5.2$\times$5.2 × EE improvement compared with random device selection.
@@ IV-C4 Data Selection
The efficiency of model training can be improved by identifying important data samples in large datasets [156]. This presents two opportunities for energy savings in FEEL. First, removing irrelevant and potentially adversarial data samples accelerates model convergence, reducing both the computation energy for local training and the communication energy for model exchange. Second, using informative data samples in local training reduces the computation overhead in each training iteration. However, evaluating the utility of data samples is challenging in FEEL, as data are kept locally at EUDs.
@@ IV-C5 Data Offloading
In FEEL applications where EUDs and MEC servers are managed by the same authority, the data privacy concerns can be much relieved, allowing data offloading to reduce on-device training overhead. The MEC server, with its more capable computational resources, is well-suited to collect data offloaded from EUDs and perform model training using the collected data [161, 162, 163]. The data offloading problem in FEEL was formulated as an MDP problem in [161], where each EUD determines whether to upload a data sample to the MEC server or use it for local training based on channel quality and battery energy state. In [162] and [163], it was proposed to offload data from EUDs to the MEC server prior to model training. The offloaded data volume and resource management were jointly optimized to minimize training latency while satisfying energy consumption constraints. In addition to vertical data offloading to the MEC server, there is also growing interest in horizontal data offloading to nearby EUDs by leveraging the device-to-device (D2D) communication links [164, 165, 166]. However, for applications with stringent data privacy requirements, offloading raw data may not be admissible. In such cases, the framework of coded FL can be adopted, where privacy-preserving coding schemes are applied to the local dataset before data offloading [167, 168, 169].
IV-D Takeaways
@@
This section introduces energy-efficient cooperative model training for edge AI, with a specific focus on FEEL as an illustrative framework. Green FEEL systems are typically realized through two techniques: compression and resource management. Compression techniques aim to reduce the amount of data communicated between the MEC server and EUDs, and the number of FLOPs required for local model training. These objectives can be achieved through methods such as quantization, sparsification, and knowledge distillation. While various approaches have been summarized in this regard, it is important to delve deeper into compression methodologies that can adapt to the inherent heterogeneity at the device, model, and data levels. This nuanced investigation is essential to align with practical usage scenarios. Moreover, privacy-enhancing techniques such as local differential privacy (LDP), can be jointly applied with compression [170]. Resource management techniques focus on efficiently utilizing the limited communication and computational resources during the training process of FEEL. This involves optimizing training parameters, such as the number of local training epochs per iteration, batch sizes, and CPU-cycle frequencies, to maximize the EE while maintaining good learning performance. Furthermore, judiciously selecting the data and devices that participate in the training procedures contributes to enhancing the EE. A crucial aspect in this area is the robust allocation of resources to handle device failures [171] within the FEEL paradigm. This practical concern warrants comprehensive investigation.
@@ V Energy-Efficient Edge Inference: Paradigms and Algorithms
Despite model training is extremly energy-demanding, most energy of DNN processing is actually consumed at the inference stage that repeats for every new observation. This section introduces three representative edge inference paradigms and the corresponding EE-oriented design methodologies. Specifially, we start from the primitive paradigms of on-device and MEC server-based inference, which respectively present high on-device computation and communication overhead. To eliminate these bottlenecks, the emerging paradigm of device-edge cooperative inference (“co-inference”) is then investigated. Fig. 5 illustrates the three typical edge inference paradigms and Fig. 6 summarizes the corresponding energy-efficient design approaches.
@@ V-A On-device Inference
On-device inference makes predictions for local observations by deploying pre-trained DNN models at EUDs. Such a paradigm fits well for EUDs with reasonable amount of computational resources and privacy-sensitive applications that prohibit data offloading. However, processing DNNs fully on-device is very challenging due to the computational resource, battery energy, and thermal constraints [172]. Previous efforts mainly target at improving the realtime inference performance [173, 174, 175] without scrutinizing the energy footprint. Next, we present three categories of techniques that can effectively trim the on-device inference energy consumption.
@@
Knowledge Distillation: KD is able to generate compressed DNN models in customized architectures according to energy budgets of EUDs. The use of different knowledge on DNN models results in different KD techniques, including response-based, feature-based, and relation-based KD [186]. Response-based KD utilizes outputs of one or more teacher DNNs to train a student model, which shows promising results on image classification and automatic speech recognition [133]. To leverage the intermediate representations learned by a teacher model, a feature-based KD method named FitNets was proposed in [187], which reduces 10×10$\times$10 × of parameters with improved inference accuracy. Moreover, considering knowledge can be better represented in the embedding space, relation-based KD exploits the mutual relations of data samples in the teacher model’s outputs [188]. It can be applied with response-based or feature-based KD, while the combination of relation-based and response-based KD achieves higher accuracy on classifying CIFAR-100 and Tiny ImageNet as observed in [188]. Recently, progressive KD was proposed in [189] that distills a large teacher model into a collection of small student models, which can be selectively activated in runtime for ensemble inference to balance the energy cost and inference accuracy.
@@
Figure 7: Input-adaptive deep neural network architectures for on-device inference.
2) Neural Architecture Search: Model compression techniques are bundled with existing DNN model architectures. Typically, DNN models are handcrafted by compositing different basic operations. For example, MobileNet [173], a streamlined architecture for mobile vision, adopts the depth-wise separable convolutions. However, hand-crafted approaches are too laborious to enumerate many possibilities. Neural architecture search (NAS), which leverages optimization methods, such as reinforcement learning and evolutionary algorithms, efficiently traverses the search space for target performance metrics. Despite primitive NAS frameworks concern more on model accuracy and computation complexity [190, 191], they have been recently generalized with energy awareness [192]. To incooporate the energy consumption, a multi-object NAS framework was developed in [193], where the energy cost of a candidate model is measured via the NVIDIA profiling tool “nvprof”. For other platforms, energy consumption can be estimated via alternative profiling methods [194] or prediction models [195]. More discussions on platform-aware NAS are deferred to Section VI-B. Besides, mixed-precision NAS, which jointly considers quantization and NAS, was investigated in [196]. It reduces the inference energy consumption on the CIFAR-100 classfication task by $20.9\%$ and $76.2\%$, compared with HAQ [180] and MobileNetV2 [197], respectively, while maintaining similar inference accuracy.
@@
3) Input-Adaptive Inference: The computation graphs of conventional DNN models are fixed. Such one-for-all designs ignore the uneven difficulty of inference data, which are suboptimal in terms of EE. Hence, various input-adaptive inference strategies [198, 199] as depicted in Fig. 7, including model selection, early termination, and layer/channel skipping, have been proposed.
@@
Model Selection: As shown in Fig. 6(a), model selection for input-adaptive inference deploys two DNN models with different computation overhead and inference accuracy. To decide which DNN to use, a lightweight binary classifier can be adopted to evaluate the input difficulty. For image classification, the two DNNs can be different pruned versions of a base DNN model, and a relabelled dataset can be obtained to train the binary classifier [200]. MobiSR [201] follows a similar idea for image super resolution, where the difficulty evaluation unit was designed with a total variation metric. It is also possible to select from more than two DNN models [202].
@@
Early Termination: Early termination employs multiple intermediate exits to decide whether to terminate the inference process at an early stage. Fig. 6(b) and Fig. 6(c) show two early termination DNN architectures, which are respectively called amortized inference [203] and branchy network [204]. In amortized inference, multiple fully functional DNNs with different computation overhead and inference accuracy are cascaded, where each input data is first processed by the small models, followed by the larger ones. The inference result of each DNN is used to compute a confidence score as an indicator for early termination. For branchy networks, early exits, which are lightweight classifiers, are inserted along the depth of a base DNN model. Since the early exits incur extra computation overhead and may even degrade the base model accuracy [205], the number and insertion locations are crucial to the EE. To address this issue, FlexDNN was developed to find an optimal insertion plan in [206] .
@@
Layer/Channel Skipping: Layer/channel skipping allows more flexible adaptation by determining which network layers/channels in a network layer should be executed according to the inference data. It can be realized by training a gating module for each candidate layer/channel [205]. The layer skipping architecture SkipNet [207] shown in Fig. 6(d) suits well for ResNets due to their robustness to layer dropping. Alternatively, BlockDrop [208] trains a policy network that outputs probabilities of binary dropping decisions for all layers under a single-step MDP formulation, which addresses both the prediction correctness and number of utilized residual blocks. To avoid overwhelming the energy savings brought by layer skipping, the gating modules/policy network should have negligible computation cost compared to the base model. Hence, it was further suggested to use a fraction of the depth in the base ResNet. Channel skipping such as CGNet [209] can be viewed as a fine-grained variant of layer skipping since the skipping policies for individual channels need to be learned. Moreover, a multi-grained approach named IDAI was proposed in [210], which supports simultaneous layer and channel skipping.
@@
Table V summarizes the computation and energy savings, as well as accuracy gain of selected input-adaptive on-device inference methods.

### v2

@@ IV-B Model and Gradient Compression
In each training iteration, the model updates to be uploaded by EUDs take the form of model weights or model gradients. Compressing these model updates can reduce communication energy consumption, as fewer bits need to be transmitted. In addition, locally training compressed models can effectively minimize on-device computation energy consumption, as it can avoid substantial floating-point/MAC and memory access operations [125, 126]. Typical compression approaches such as quantization, sparsification, and knowledge distillation are often employed to achieve energy-efficient FEEL.
@@ IV-B1 Quantization
In FEEL, quantization employs low bitwidths to represent DNN model parameters on-device. This technique can also be applied to model gradients prior to transmission. Given that quantization is an irreversible process, it is crucial to balance achieving maximum EE with maintaining satisfactory learning performance.
@@ IV-B3 Knowledge Distillation
A well-known KD-based FL algorithm called FedMD [144] addresses this issue by allowing EUDs to have heterogeneous model architectures based on their resource constraints. To facilitate cooperative training, logits computed from a public dataset using local models are uploaded to and aggregated at the MEC server. These aggregated logits are then redistributed to EUDs for local training using TL and KD. Since logits are of low dimension, the communication overhead of FedMD is significantly reduced compared to FedAvg. Empirical results in [145] provide further insights into key parameters of KD-based FEEL, such as the size and distribution of the public dataset, as well as the logit aggregation strategy. Another approach, FedKD [146], proposes maintaining a teacher and a student model at each EUD that can learn from each other, with only the approximated student model gradients, which have much smaller sizes, being uploaded. Complexity analysis demonstrates that FedKD significantly reduces communication costs with a marginal increase in local computation energy consumption. Besides, a selective knowledge sharing mechanism was developed for KD-based FL in [147], which identifies accurate and precise knowledge respectively from local and ensemble predictions over the learning process. KD can also be implemented at the MEC server to simplify computation at EUDs. For example, in Mix2FLD [148], each EUD uploads a set of mixed-up data samples along with the logits. The logit-to-model conversion is performed by running a KD algorithm at the MEC server, after which the updated global model is downloaded by the EUDs. Mix2FLD demonstrates faster convergence and higher model accuracy, implying much lower energy consumption at the EUDs.
@@ IV-C1 Local Training Adaptation
Since most computations in FEEL are disseminated to EUDs, the approach to local training in each iteration has direct impacts on learning performance, as well as time and energy efficiency. Intuitively, better training of local models in each iteration is beneficial of reducing the required number of iterations to achieve a learning performance target [123, 149, 150], but it may increase the on-device processing and computation energy consumption per iteration. However, a smaller number of iterations is advantageous for saving communication energy throughout the training process. Therefore, local training in FEEL should be meticulously adapted to find the optimal balance.
In [151], local training adaptation was first studied for FEEL, where a convergence bound was derived considering arbitrary local training epochs in each global iteration. This bound was utilized to determine the number of local epochs that minimizes the loss function while adhering to resource constraints, such as energy and monetary costs. The batch size of local SGD also plays a significant role, as a larger batch size reduces gradient variance but increases computation overhead [152]. To conserve battery energy by minimizing device idling, an efficient algorithm was developed in [153] to dynamically determine batch sizes and learning rates for heterogeneous EUDs. Additionally, scaling the on-device processors to appropriate speeds can improve the EE of local training. For instance, in [154], the total system cost of FEEL defined as a weighted sum of training time and energy consumption was minimized using an experience-driven central processing unit (CPU)-cycle frequency control algorithm. Compared to using a static CPU-cycle frequency, this approach achieves over $4\%$ energy savings with a substantial reduction in training time.
@@ IV-C2 Joint Computation-Communication Control
Recent studies have also incorporated state-of-the-art computing architectures [157] and air interface technologies [158]. In [157], each EUD was assumed to be equipped with a CPU-GPU heterogeneous computing platform. Consequently, the joint optimization of communication resource management, CPU-GPU workload partitioning, and processor frequency scaling was performed to minimize total energy consumption. Experimental results demonstrated the effectiveness of joint resource management in improving EE compared to separate management of heterogeneous resources. Attention was also given to the intelligent reflecting surface (IRS) technology, which utilizes passive elements such as diodes and phase shifters to customize wireless environments in an energy-efficient manner. In [158], an IRS was proposed to be deployed in FEEL systems to provide favorable wireless channels, with the phase shifts at the IRS and the local CPU speed being judiciously designed. Simulation results showcased speedy convergence and significant energy savings, especially with a large-scale IRS.
@@ IV-C3 Device Selection
In [161], a device selection problem was tackled by a hybrid branch-and-bound algorithm in each training iteration to minimize the energy consumption of participating EUDs in a TDMA-based FEEL system. Additionally, the trade-off between maximizing the number of selected EUDs and minimizing their total energy consumption was investigated in [162], leading to the development of an energy- and latency-aware device selection algorithm. Both [161] and [162] demonstrate the effectiveness of device selection in achieving energy-efficient FEEL. To ensure unbiased model aggregation with a subset of selected devices, probabilistic device scheduling frameworks were developed in [163, 164].
The aforementioned studies focus on optimizing device selection independently in each training iteration. However, all decisions made throughout the training process can have a significant impact on the learning outcome. This motivates the design of device selection strategies from a long-term perspective [165, 166, 167, 168]. In particular, recognizing that the number of participating EUDs in later training rounds has a more profound impact on learning performance, device selection and bandwidth allocation were jointly optimized in [165] using Lyapunov optimization, subject to energy constraints at each EUD. Experimental results demonstrate the benefits of this approach over methods that myopically minimize energy consumption in each round. This work was further extended in [166] to address the device heterogeneity issue by exploiting sequential transmission of local model updates, where early available local model updates can be uploaded to the MEC server using more bandwidth. Unlike [165] and [166], which use the average number of participating EUDs as a surrogate for learning performance, [167] developed an online data-driven method to estimate learning performance using a validation dataset. Based on this estimation, the selected devices and resource management were jointly optimized to minimize the system energy consumption while respecting latency and accuracy requirements. In [168], a lightweight AutoFL framework was developed based on reinforcement learning to select the best EUDs in each training round, achieving 5.2×5.2$\times$5.2 × EE improvement compared with random device selection.
@@ IV-C4 Data Selection
The efficiency of model training can be improved by identifying important data samples in large datasets [169]. This presents two opportunities for energy savings in FEEL. First, removing irrelevant and potentially adversarial data samples accelerates model convergence, reducing both the computation energy for local training and the communication energy for model exchanging. Second, using only the most informative data samples in local training reduces the computation overhead in each training iteration. However, evaluating the utility of data samples is challenging in FEEL, as data are kept locally at EUDs.
@@ IV-C5 Data Offloading
In FEEL applications where EUDs and MEC servers are managed by the same authority, the data privacy concerns can be much relieved, allowing data offloading to reduce on-device training overhead. The MEC server, with its more capable computational resources, is well-suited to collect data offloaded from EUDs (i.e., vertical data offloading) and perform model training using the collected data [174, 175, 176]. The data offloading problem in FEEL was formulated as an MDP problem in [174], where each EUD determines whether to upload a data sample to the MEC server or use it for local training based on channel quality and battery energy state. In [175] and [176], it was proposed to offload data from EUDs to the MEC server prior to model training. The offloaded data volume and resource management were jointly optimized to minimize training latency while satisfying energy consumption constraints. In addition to vertical data offloading to the MEC server, there is also growing interest in horizontal data offloading to nearby EUDs by leveraging the device-to-device (D2D) communication links [177, 178, 179]. However, for applications with stringent data privacy requirements, offloading raw data may not be admissible. In such cases, the framework of coded FL can be adopted, where privacy-preserving coding schemes are applied to the local dataset before data offloading [180, 181, 182].
IV-D EE-oriented Designs for Other Cooperative Edge Learning Frameworks
In this subsection, we introduce two alternative cooperative edge learning frameworks, including DEEL and H-FEEL, and elaborate several of their EE-oriented designs.
IV-D1 DEEL
Contrary to FEEL systems that employ an MEC server as the PS, EUDs in DEEL systems share local model updates with their one-hop neighboring EUDs in each training iteration, which is more resilient to node failures. Model aggregation is performed at individual EUDs. Typically, an EUD generates a new local model by mixing its own model and those from neighboring EUDs with appropriate weighting factors [121]. Despite the training algorithm of DEEL is different from that of FEEL, most design principles of energy-efficient FEEL are applicable for achieving energy-efficient DEEL. For example, a DEEL mechanism where EUDs exchange quantized model parameters was developed in [183]. Experimental results on a beat signal dataset show that the model accuracy just slightly degrades even with 1-bit quantization. Since communications in DEEL take place between neighboring EUDs, similar as FEEL, wireless links in unfavorable conditions may prolong the learning latency and result in waste of transmission energy. Therefore, link selection strategies should be optimized apart from communication and computational resource management [184, 185]. To reduce both the communication and computation energy consumption, joint model pruning and network topology construction for DEEL was investigated in [186], where each EUD trains a differently pruned sub-model and the network topology is adapted according to data heterogeneity and link speeds.
IV-D2 H-FEEL
In H-FEEL systems, EUDs are partitioned into multiple clusters and EUDs in a cluster are associated with an MEC server. The MEC servers are coordinated by a cloud server [122]. Hence, model aggregation of H-FEEL consists of two levels, namely the edge- and cloud-level aggregation. Such a cooperative edge model training paradigm extends the network coverage to engage many EUDs, thus improving the learning performance. As the operations within each cluster of H-FEEL are largely the same as those of FEEL, it can be expected that the design approaches discussed in Section IV-B and IV-C remain effective [187, 188]. For instance, in [187], convergence performance of H-FEEL with quantized gradient transmission and imperfect CSI was analyzed, which unveils the relationship between quantization accuracy and transmission error. Adaptive model pruning was adopted for H-FEEL in [188] to minimize the communication and computation overhead. To determine the prune ratios of EUDs and bandwidth allocation in each cluster, a minimization problem on a convergence bound was optimally solved. Compared to FEEL, device-edge association schemes in H-FEEL have profound impact on the learning performance and cost, which thus become a new frontier to be studied [189, 190]. Specifically, communication-computation resource allocation and device-edge association was jointly optimized for H-FEEL to minimize the global energy and latency cost. Noticing that the imbalanced data distribution across clients is a main cause of slow convergence and degraded accuracy, device-edge association was obtained by minimizing the divergence between the federated and hypothetical central models in [190]. As such, the communication and computation energy consumption can be reduced with less required training rounds.
IV-E Takeaways
@@
This section introduces energy-efficient cooperative model training for edge AI, with a specific focus on FEEL as an illustrative framework. Green FEEL systems are typically realized through two techniques: compression and resource management. Compression techniques aim to reduce the amount of data communicated between the MEC server and EUDs, and the number of FLOPs required for local model training. These objectives can be achieved through methods such as quantization, sparsification, and knowledge distillation. While various approaches have been summarized in this regard, it is important to delve deeper into compression methodologies that can adapt to the inherent heterogeneity at the device, model, and data levels. This nuanced investigation is essential to align with practical usage scenarios. Moreover, privacy-enhancing techniques such as local differential privacy (LDP), can be jointly applied with compression [191]. Resource management techniques focus on efficiently utilizing the limited communication and computational resources during the training process of FEEL. This involves optimizing training parameters, such as the number of local training epochs per iteration, batch sizes, and CPU-cycle frequencies, to maximize the EE while maintaining good learning performance. Furthermore, judiciously selecting the data and devices that participate in the training procedures contributes to enhancing the EE. A highly important aspect in this area is the robust allocation of resources to handle device failures that might occasionally happen due to energy shortage [192] within the FEEL paradigm. Besides, EE analysis for large-scale FEEL systems is instructive for real-world deployments, for which, stochastic geometry theory can provide a powerful framework [193]. These practical concerns warrant comprehensive investigations. Moreover, although the EE-oriented design approaches of FEEL can be leveraged for other cooperative edge learning frameworks, attentions on their unique operations should be given.
@@ V Energy-Efficient Edge Inference: Paradigms and Algorithms
Despite model training is extremely energy-demanding, most energy of DNN processing is actually consumed at the inference stage that repeats for every new observation. This section introduces three representative edge inference paradigms [14] and the corresponding EE-oriented design methodologies. Specifically, we start from the primitive paradigms of on-device and MEC server-based inference, which respectively present high on-device computation and communication overhead. To eliminate these bottlenecks, the emerging paradigm of device-edge cooperative inference (“co-inference”) is then investigated. Fig. 6 illustrates the three typical edge inference paradigms and Fig. 7 summarizes the corresponding energy-efficient design approaches.
@@ V-A On-device Inference
On-device inference makes predictions for local observations by deploying pre-trained DNN models at EUDs. Such a paradigm fits well for EUDs with a reasonable amount of computational resources and privacy-sensitive applications that prohibit data offloading. However, processing DNNs fully on-device is very challenging due to the computational resource, battery energy, and thermal constraints [194]. Previous efforts mainly target at improving the real-time inference performance [195, 196, 197] without scrutinizing the energy footprint. Next, we present three categories of techniques that can effectively trim the on-device inference energy consumption.
@@
Knowledge Distillation: KD is able to generate compressed DNN models in customized architectures according to energy budgets of EUDs. The use of different knowledge on DNN models results in different KD techniques, including response-based, feature-based, and relation-based KD, as summarized in [208]. Specifically, response-based KD utilizes logits of one or more teacher DNNs to train a student model, which shows promising results on image classification and automatic speech recognition [143]. To leverage the intermediate representations learned by a teacher model, a feature-based KD method named FitNets was proposed in [209], which reduces 10×10$\times$10 × of parameters compared to a larger teacher DNN while with improved inference accuracy. Moreover, considering knowledge can be better represented in the embedding space, relation-based KD exploits the mutual relations of data samples in the teacher model’s outputs [210]. It can be applied with response-based or feature-based KD, while the combination of relation-based and response-based KD achieves higher accuracy on classifying CIFAR-100 and Tiny ImageNet as observed in [210]. Recently, progressive KD was proposed in [211] that distills a large teacher model into a collection of small student models, which can be selectively activated in runtime for ensemble inference to balance the energy cost and inference accuracy.
2) Neural Architecture Search: Model compression techniques are bundled with existing DNN model architectures. Typically, DNN models are handcrafted by compositing different basic operations. For example, MobileNet [195], a streamlined architecture for mobile vision, adopts the depth-wise separable convolutions. However, handcrafted approaches are too laborious to enumerate many possibilities. Neural architecture search (NAS), which leverages optimization methods, such as reinforcement learning and evolutionary algorithms [212], efficiently traverses the search space for target performance metrics. Despite primitive NAS frameworks were concerned more on model accuracy and computation complexity [212], they have been later generalized with energy awareness [213, 214]. To incorporate the energy consumption, a multi-object NAS framework was developed in [215], where the energy cost of a candidate model is measured via the NVIDIA profiling tool “nvprof”. For other platforms, energy consumption can be estimated via alternative profiling methods (e.g., by measuring the current and voltage of RPi) [216] or prediction models (e.g., random forest) [217]. More discussions on platform-aware NAS are deferred to Section VI-B. Besides, mixed-precision NAS, which jointly considers quantization and NAS, was investigated in [218]. It reduces the inference energy consumption on the CIFAR-100 classification task by $20.9\%$ and $76.2\%$, compared with HAQ [202] and MobileNetV2 [219], respectively, while maintaining similar inference accuracy.
@@
Figure 8: Input-adaptive deep neural network architectures for on-device inference, including (a) model selection [220], (b) cascaded inference [221], (c) branchy network [222], and (d) layer skipping [223].
@@
3) Input-Adaptive Inference: The computation graphs of conventional DNN models are fixed. Such one-for-all designs ignore the uneven difficulty of inference data, which are suboptimal in terms of EE. Hence, various input-adaptive inference strategies [224, 225, 226] with examples depicted in Fig. 8, including model selection, early termination, and layer/channel skipping, have been proposed.
@@
Model Selection: As shown in Fig. 8(a), model selection for input-adaptive inference deploys two DNN models with different computation overhead and inference accuracy. To decide which DNN to use, a lightweight binary classifier can be adopted to evaluate the input difficulty. For image classification, the two DNNs can be different pruned versions of a base DNN model, and a relabelled dataset can be obtained to train the binary classifier [220]. MobiSR [227] follows a similar idea for image super resolution, where the upscaling difficulty evaluation unit was designed with a total variation metric. It is also possible to select from more than two DNN models [228].
@@
Early Termination: Early termination employs multiple intermediate exits to decide whether to terminate the inference process at an early stage. Fig. 8(b) and Fig. 8(c) show two early termination DNN architectures, which are respectively called cascaded inference [221] and branchy network [222]. In cascaded inference, multiple fully functional DNNs with different computation overhead and inference accuracy are concatenated, where each input data is first processed by the small models, followed by the larger ones. The inference result of each DNN is used to compute a metric (e.g., confidence score) as an indicator for early termination. For branchy networks, early exits, which are lightweight classifiers, are inserted along the depth of a base DNN model. Since the early exits incur extra computation overhead and may even degrade the base model accuracy [226], the number and insertion locations are crucial to the EE. To address this issue, FlexDNN was developed to find an optimal insertion plan in [229] .
@@
Layer/Channel Skipping: Layer/channel skipping allows more flexible adaptation by determining which network layers/channels in a network layer should be executed according to the inference data. It can be realized by training a gating module for each candidate layer/channel [226]. For example, the layer skipping architecture SkipNet [223] shown in Fig. 8(d) suits well for ResNets due to its robustness to layer dropping. Alternatively, BlockDrop [230] trains a policy network that outputs probabilities of binary dropping decisions for all layers under a single-step MDP formulation, which addresses both the prediction correctness and number of utilized residual blocks. To avoid overwhelming the energy savings brought by layer skipping, the gating modules/policy network should have negligible computation cost compared to the base model. Hence, it was further suggested in [230] to use a fraction of the depth in the base ResNet as the policy network architecture. Channel skipping such as CGNet [231] can be viewed as a fine-grained variant of layer skipping since the skipping policies for individual channels need to be learned. Moreover, a multi-grained approach named IADI was proposed in [232], which supports simultaneous layer and channel skipping.
@@
Table V summarizes the computation and energy savings, as well as accuracy gain of selected input-adaptive on-device inference methods. Interested readers may refer to [226] for more discussions on the taxonomy and architectures of input-adaptive inference.
V-B MEC Server-based Inference